Hook Industries' capital structure consists solely of debt and common equity. It can issue debt at rd = 12%, and its common stock currently pays a $2.25 dividend per share (D0 = $2.25). The stock's price is currently $29.50, its dividend is expected to grow at a constant rate of 9% per year, its tax rate is 35%, and its WACC is 13.90%. What percentage of the company's capital structure consists of debt? Round your answer to two decimal places.
Answer #1

WACC = Weight of debt * Pretax cost of debt * (1 - Tax) + Weight of Equity * Cost of Equity

Cost of Equity can be calculated using dividend discount model.

This question requires application of constant growth dividend discount model according to which: Po - Divi T-9 Po = Price of

Div1 = $2.25 * (1 + 9%) = $2.4525

29.50 = 2.4525 r - 0,09

r - 0.09 = 0.0831

Cost of Equity, r = 17.31%

Weight of Debt + Weight of Equity = 1

Weight of Equity = 1 - Weight of Debt

13.90% = Weight of debt * 12% * (1 - 35%) + Weight of Equity * 17.31%

13.90% = Weight of debt * 12% * (1 - 35%) + (1 - Weight of Debt) * 17.31%

13.90% = Weight of debt * 7.80% + 17.31% - Weight of Debt * 17.31%

13.90% - 17.31% = - 9.51% * Weight of Debt

-3.41% = - 9.51% * Weight of Debt

Weight of debt = 35.88%